### v2

# Genomic Next-Token Predictors are In-Context Learners

Abstract
In-context learning (ICL) – the capacity of a model to infer and apply abstract patterns from examples provided within its input – has been extensively studied in large language models trained for next-token prediction on human text. In fact, prior work often attributes this emergent behavior to distinctive statistical properties in human language. This raises a fundamental question: can ICL arise organically in other sequence domains purely through large-scale predictive training?
To explore this, we turn to genomic sequences, an alternative symbolic domain rich in statistical structure. Specifically, we study the Evo2 genomic model, trained predominantly on next-nucleotide (A/T/C/G) prediction, at a scale comparable to mid-sized LLMs. We develop a controlled experimental framework comprising symbolic reasoning tasks instantiated in both linguistic and genomic forms, enabling direct comparison of ICL across genomic and linguistic models. Our results show that genomic models, like their linguistic counterparts, exhibit log-linear gains in pattern induction as the number of in-context demonstrations increases. To the best of our knowledge, this is the first evidence of organically emergent ICL in genomic sequences, supporting the hypothesis that ICL arises as a consequence of large-scale predictive modeling over rich data. These findings extend emergent meta-learning beyond language, pointing toward a unified, modality-agnostic view of in-context learning.
1 Introduction
footnotetext: LLMs were used during the conception, implementation, and refinement of this paper. We assume full responsibility for all content.
Scaling Large Language Models (LLMs) has revealed an unexpected and powerful capacity: in-context learning (ICL) (radford2018improving; brown2020language), the ability to infer and apply abstract patterns purely from examples contained within their input. Unlike conventional adaptation, which relies on gradient updates, LLMs internalize mechanisms for flexible pattern induction (olsson2022context), allowing them to perform few-shot generalization and analogical reasoning (srivastava2023beyond), without explicit parameter updates. This behavior, which arises organically from large-scale “next-token” prediction on human language, has been interpreted as emergent forms of meta-learning (learning to learn) within a fixed parametric system.
Almost all prominent evidence of emergent ICL so far (srivastava2023beyond, inter alia) comes from training on human language (e.g., English). This raises a fundamental question: is there something inherently special about human language that enables ICL to emerge? One can take two different stances on this:
  • H1: Human language possesses distinctive distributional properties such as parallelism or compositionality (chen2024parallel; hahn2023theory) that uniquely nurture ICL.
  • H2: Alternatively, perhaps natural language is merely a convenient substrate – having ample data and being naturally interpretable – and ICL could just as well arise in any sufficiently structured data domain, provided the model is large enough and trained to compress predictive patterns.
While many works have examined H1, the veracity of H2 remains far less understood. To this end, we turn to a radically different substrate: the genome. Genomic sequences can be viewed as a form of natural language that has evolved through nature. Like human language, they comprise complex symbol sequences that exhibit rich statistical regularities – motifs, repeats, dependencies (benegas2025genomic) that could, in principle, support pattern induction within context.
Testing for the existence of ICL requires large-scale models. Historically, genomic models were small and specialized, optimized for narrow biological tasks such as motif discovery or structure prediction. The advent of deep biological models (Ji2020DNABERTPB) has expanded their scale, paralleling those of early LLMs. Most recently, the Evo2 series (evo2report) marks a major step forward: a large-scale model trained solely on ‘next-nucleotide’ prediction. With training scale comparable to mid-sized LLMs (e.g., Qwen3-14B (yang2025qwen3technicalreport)), these genomic models finally make it possible to systematically compare genomic and linguistic representations under large-scale autoregressive training. If ICL arises primarily from scale and predictive compression (H2), then large-scale genomic models such as Evo2 must already exhibit it under our noses. Like LLMs, they are trained for next-token prediction without any explicit meta-learning signal. Yet, despite extensive analysis of Evo2’s biological capabilities (e.g., protein fitness prediction, variant effect estimation), its potential for emergent few-shot or analogical reasoning within context remains unexplored.
Figure 1: We design parallel symbolic reasoning tasks that allow direct comparison of ICL behavior across modalities (linguistic and genomic). 1 Few-shot bitstring program-synthesis tasks (e.g., identity, NOT, majority, reverse) require models to infer mappings from examples. 2 Each task is rendered in two modality-specific encodings: genomic (bitstrings mapped to random nucleotides A/T/C/G) and linguistic (bitstrings mapped to random digits), preserving abstract structure but differing in surface form. 3 Both genomic (Evo2) and linguistic (Qwen3) models receive $k$-shot demonstrations and are greedily decoded to compute exact-match accuracy. 4 Both models show log-linear accuracy gains with more demonstrations.
The second challenge in evaluating H2 lies in developing a framework to detect and quantify ICL in the genomics domain where “tasks” lack natural human interpretability. Moreover, the task design must permit systematic comparison between ICL behaviors in human language and genomic modalities. To investigate this, in §3 we introduce a controlled experimental framework (Fig. 1) for evaluating ICL across linguistic and genomic modalities. As illustrated in 12, we design a suite of symbolic bitstring program-synthesis tasks and render them in two parallel encodings: one using genomic alphabets (nucleotides: A/T/C/G) and one using linguistic (characters). We then evaluate both model families – Evo2 for genomic sequences and Qwen3 for linguistic text – under matched few-shot prompting conditions ( 3). This parallel encoding enables direct, apples-to-apples comparison of ICL scaling trends across genomic and linguistic models.
Our experiment’s results (§4) reveal striking parallels: both linguistic and genomic models exhibit log-linear improvements in pattern induction and extrapolation as the number of demonstrations grows. To our knowledge, this is the first evidence of organically emergent ICL in genomic sequences. These findings suggest H2, that is, ICL is not an artifact that is specific to human language, but a broader consequence of compression and predictive modeling in pattern-rich sequence spaces. By demonstrating ICL in a non-linguistic biological domain, we extend the scope of emergent meta-learning beyond natural language, pointing toward a unified view of context-adaptive computation that spans different modalities of data.
We further perform analyses to compare the types of tasks at which genomic and linguistic models excel. Our results reveal distinct patterns: genomic models learn faster with more demonstrations, while linguistic models gain more from scale (§4.2); genomic models perform best on direct copying or local transformations, whereas linguistic models excel on tasks requiring single-bit dependencies or global summary statistics (§4.3, §4.4). We note the caveat that these findings are specific to our experimental conditions (e.g., our choice of reasoning tasks) and may not be a holistic statement about ICL in genomic and linguistic models.
Our Contributions. (a) We present a controlled experimental setup for detecting and measuring ICL across linguistic and genomic modalities. Our framework defines a suite of symbolic bitstring reasoning tasks that can be rendered in both representations, enabling direct, apples-to-apples comparison of ICL behavior. (b) Using the Evo2 family of large-scale genomic models, we demonstrate for the first time that models trained solely on nucleotide sequences exhibit clear in-context learning, mirroring the scaling trends of Qwen3 language models. (c) Our findings challenge the idea that ICL depends on linguistic structure. Instead, they point to ICL as a modality-agnostic outcome of large-scale next-token prediction over pattern-rich data, reflecting general principles of compression and contextual inference. Together, these results extend the scope of emergent meta-learning beyond language, toward a unified understanding of context-adaptive computation in artificial and biological systems alike.
2 Related Work and Broader Context
Why does ICL remain interesting? ICL remains a striking and conceptually rich phenomenon: it emerges without task-specific training yet enables rapid, in-situ adaptation. It comes closest to long-standing ambitions in classical AI—such as case-based (aamodt1994case; lancaster1987problem; ross1984remindings; ellman1989explanation) and analogical reasoning (hofstadter2001analogy; gentner2017analogy; holyoak2001place) where agents solve new problems by reusing solutions from prior (analogous) examples.
Beyond its theoretical appeal, ICL has become a practical workhorse in recent advances, e.g., synthetic data generation using pretrained models (wang2023selfinstruct; shao2023synthetic; tunstall2023zephyr) or reasoning (wei2022chain; nye2021show; yao2022react; yasunaga2023large). In short, ICL is both a window into emergent learning dynamics and a cornerstone for building adaptive, interpretable systems.
Emergent-ICL $\neq$ Meta-ICL: A parallel line of research studies models trained with an explicit ICL objective – learning to predict over collections of input–output pairs $(x,f(x))$. For example, regression models that are trained explicitly to infer an underlying rule (e.g., linear, polynomial, or sinusoidal) from few-shot examples (garg2022can; li2023transformers; raventos2024pretraining; nejjar2024context). This setup mirrors meta-learning, where a model is explicitly “trained to learn” (finn2017model; bertinetto2018meta; zintgraf2019fast). We refer to such systems as Meta-ICL (panwar2023context; min2022metaicl; wu2022continued; kirsch2022general; zhang2023incontext), in contrast to emergent ICL in large pretrained models, which arises organically from next-token prediction over natural data.
Meta-ICL results demonstrate that transformers can be optimized to perform meta-learning but fail to explain why emergent ICL appears in models never trained for it. Critically, Meta-ICL generalizes poorly outside its training domain (naim2024re). For example, a Meta-ICL model trained on linear and cosine functions fails on their composition (yadlowsky2023pretraining)—and does not capture higher-order abilities like in-context reasoning (wei2022chain; kojima2022large) or adaptive self-refinement (madaan2023self) – unless explicitly trained for them. In contrast, emergent ICL in LMs generalizes broadly and flexibly across diverse problem types (srivastava2023beyond). These behavioral and mechanistic differences support our working assumption that Meta-ICL and emergent ICL are fundamentally distinct phenomena (shen2024icl_vs_gd; mishra2025ia2alignmenticlactivations). Throughout, our focus in this work is on the latter – the organic emergence of in-context learning in large pretrained models.
Are there any prior results on emergent ICL in genomic models? Most genomic models to date are encoders (Ji2020DNABERTPB; Dallatorre2024NucleotideTB), which preclude autoregressive generation and the study of ICL in the standard sense. Among the few autoregressive genomic models, most remain small, e.g., GENA-LM (300M) (Fishman2024GENALMAF) and scGPT (53M) (cui2024scgpt). The recently-released Evo2 series (evo2report) includes models up to 40B parameters, comparable in scale to modern LLMs. This model family provides the first viable testbed for probing emergent ICL, which we study here. A related effort, HyenaDNA (nguyen2023hyenadna), trains on million-token genomic sequences and reports ICL-like behaviors; however, these involve meta-ICL mechanisms (e.g., soft prompting and instruction fine-tuning; see their Sec. 4.3), rather than organically emergent ICL. An additional avenue of research demonstrates that genomic models can perform a form of ICL by copying prior tokens in-context for prediction (kantroo2025incontextlearningdistortrelationship) – akin to how induction heads function (olsson2022incontextlearninginductionheads). This, however, doesn’t demonstrate any ability to infer latent functions or apply a novel rule in-context.
Are there instances of emergent ICL in other non-linguistic modalities? While transformers have been extended to many modalities beyond language, genuine cases of emergent ICL arising purely from large-scale pretraining remain rare. In neuro-signal modeling, for instance, EEG-GPT (kim2024eeg) is trained with explicit in-context demonstrations, making it an instance of Meta-ICL rather than emergent adaptation. Similarly, multimodal models such as Flamingo (alayrac2022flamingo), BLIP/BLIP-2 (li2022blip; li2023blip2), and Emu (sun2023generative; sun2024generative) exhibit in-context behavior only because they are explicitly trained to process demonstration patterns – again, meta-learned rather than emergent. To our knowledge, the only claimed instance of emergent ICL outside language is in vision:  bai2024sequential show that transformers pretrained on natural visual sequences can infer analogical and compositional tasks without explicit supervision. The weakness of this result is that their ICL prompts consist of long sequences of contiguous frames (bai2024sequential, Fig 10), which effectively reduces the task to next-frame prediction, rather than true ICL that requires inferring compositions of multiple demonstrations.
3 Experimental Framework for Cross-domain In-context Learning
This section presents our framework for evaluating ICL across linguistic and genomic models. We outline the experimental desiderata (§3.1), setup (§3.2), model selection (§3.3), and evaluation protocol (§3.4). We additionally release code for replicating these experiments on Github.
3.1 Experimental Desiderata
Desideratum 1: Cross-domain comparability. Our experimental framework requires that each task be performable by both language and genomic models. Thus, every task must be representable in both linguistic and nucleotide alphabets. This constraint excludes existing benchmarks that rely on domain-specific semantics – such as language reasoning datasets (e.g., BIG-Bench, StrategyQA, GSM8K (srivastava2023beyond; mor2021strategyqa; cobbe2021gsm8k)) or biological tasks (e.g. variant effect prediction, exon identification  (evo2report)). While one could, in principle, translate domain-specific tasks into an alternate alphabet (e.g., mapping language tokens to base sequences via quaternary encoding), doing so inherently biases the evaluation: the source domain retains advantages, while the target domain must operate on representations that are unnatural to it. As a result, such cross-domain encodings confound the comparison, reflecting representational translation artifacts rather than genuine differences in ICL behavior.
Desideratum 2: Limited vocabulary. Since genomic models operate on only four nucleotides (A, T, C, G), tasks must be expressible within an equally compact alphabet. This rules out the existing symbolic reasoning and analogy benchmarks (hodel2023response; lewis2024evaluating; webb2025evidence), which rely on richer vocabularies (e.g., shapes, colors, or linguistic tokens with explicit semantic roles). Such tasks cannot be faithfully represented in a four-token regime without introducing artifacts or structural loss. Accordingly, our evaluation focuses on tasks that retain abstract reasoning structure while remaining compatible with the low-vocabulary symbolic space shared across linguistic and genomic models.
3.2 Experimental Setup
Notation: We formally define ICL as follows. Let $S$ be the input domain and $O$ the output domain, and let a task be a latent deterministic function $f:S\to O$. A pretrained autoregressive model $M$ performs ICL by conditioning on an ordered $n$-shot demonstration set $E=\big((x_{1},f(x_{1})),\ldots,(x_{n},f(x_{n}))\big)$, where $x_{i}\in S$, and then receiving a held-out query $x\in S\setminus\{x_{1},\dots,x_{n}\}$. Conditioned on $(E,x)$, the model produces a prediction $\hat{y}=M(E,x)$, which we compare to the ground-truth $f(x)$. Then the single-trial success is $\mathbbm{1}\big[\hat{y}=f(x)\big]$.
Task Formulation: in-context program induction over abstract symbols. Given our desiderata (§3.1), we design symbolic induction tasks that require a small lexicon. Each task requires a model to infer a hidden transformation rule from a few input–output examples and apply it to a new query. We refer to this process as program synthesis in-context. This setup isolates in-context learning ability by removing the influence of pretrained world knowledge. If the function family $f$ and symbol space $S$ are chosen carefully, the evaluation is far from the model’s pretraining distribution. Similar formulations have been used in prior work on compositional reasoning (gpt3fewshot), analogical reasoning via Raven’s Progressive Matrices (Raven1962ManualFR; webb2023emergent), and more recently in ARC-AGI (chollet2019measure).
Bitstring domain and function space. To maintain compatibility across genomic and linguistic models, we represent all symbols as bitstrings of length $k$, i.e., $S=\{0,1\}^{k}$. Genomic models operate over four nucleotides (A, C, G, T), allowing two tokens (e.g., A/C) to encode 0/1 and reserving the others (G/T) as delimiters or separators. Bitstrings are also naturally supported by linguistic models since most tokenizers represent digits as single tokens, ensuring parity in symbol granularity. This design minimizes tokenization confounds while providing a uniform symbolic substrate for comparing both domains. See Fig. 1 for examples. Despite their simplicity, bitstrings support a wide range of composable operations – from basic (identity, constant) to logical (AND, OR), positional (shift, rotation, reversal), and aggregate (majority, parity). This expressivity makes them a convenient and extensible testbed for cross-domain in-context learning evaluation.
We set the symbol space to $S=\{0,1\}^{8}$, corresponding to all 8-bit binary strings. This choice balances expressivity (256 unique bitstrings) with manageable prompt length. To define the task space, we construct a set of 100 transformation functions $F\subseteq\{f\mid f:S\to S\}$, where each function maps one bitstring to another according to a deterministic rule. Each $f\in F$ is either a single primitive operation or a composition of two primitives, allowing for a controlled yet diverse range of symbolic transformations.
Primitive operations. We build $F$ from a library of 30 fundamental primitives (listed in Appendix A.1) spanning six functional categories: (1) Bitwise transformations (e.g., flipping individual bits), (2) Structural rearrangements (rotations, shifts, reversals, swaps), (3) Positional masking and selection (preserving or zeroing specific bit positions), (4) Pattern detection (detecting alternating or palindromic patterns), (5) Aggregation functions (majority/minority, parity), and (6) Trivial mappings (identity, constant outputs). This taxonomy ensures coverage of both low-level logical operations and higher-order compositional structure.
Automatic generation and validation. To construct the final function set, we use GPT-5-Codex to automatically generate 100 unique transformations by composing primitives into valid Python programs. Each candidate function is verified programmatically for logical distinctness: no two programs $f,g\in F$ produce identical outputs for all inputs ($\nexists f,g$ such that $f(x)=g(x)$ for all $x\in S$). All functions undergo manual inspection to confirm correctness, diversity, and adherence to the intended categories. The full list of transformations and corresponding source code is provided in Appendix A.2.
Prompt encoding. Prompts are encoded using a unified symbolic scheme to enable evaluation across linguistic and genomic models. For linguistic models, bits (0,1{0,1}) are mapped to two random digits (0–9); for genomic models, to random nucleotides ($\mathrm{A},\mathrm{T},\mathrm{C},\mathrm{G}$}). The separator token ($\to$) is omitted, and in-context examples are separated by a randomly chosen unused token distinct from those representing 0 and 1. (See Fig. 1 for examples.) All mappings are randomized per trial to avoid memorization or positional bias.
3.3 Model Families and Selection Rationale
For fair cross-domain comparison, we use two representative model families: Qwen3 for human language and Evo2 for genomics. The rationale for this selection is as follows: (a) Parameter scaling: Both families span multiple orders of magnitude in parameter count, enabling systematic scaling analysis of ICL ability. The Qwen3 series ranges from 0.6B to 14B parameters, while Evo2 includes 1B, 7B, and 40B models (yang2025qwen3technicalreport; evo2report). This parallel scaling structure facilitates consistent measurement of how ICL performance evolves with model size across linguistic and genomic modalities. (b) Compute matching: The largest models in each family are trained with comparable total compute, offering an opportunity for an approximately compute-matched cross-domain comparison. Using the standard $6ND$ estimate (kaplan2020scaling), Qwen3-14B-Base is trained with about $3.2\times 10^{24}$ FLOPs, while Evo2-40B is trained with $2.25\times 10^{24}$ FLOPs (yang2025qwen3technicalreport; evo2report), making Evo2 uniquely well-suited for comparison with Qwen3 at scale. (c) Availability of base models: The Qwen3 family releases base (pre-instruction-tuned) models at all scales up to 14B parameters, enabling direct evaluation of the intrinsic inductive reasoning ability of pure next-token predictors, without instruction-tuning artifacts. The Evo2 base models have no additional instruction tuning applied – the only nuance is that Evo2 1B is trained at a context length of 8192 nucleotides, whereas the 7B/40B are extended to a context length of one million. (d) Licensing: All models are released under Apache 2.0 (yang2025qwen3technicalreport; arcinstitute_evo2_1b_base_hf_2025; arcinstitute_evo2_7b_hf_2025; arcinstitute_evo2_40b_hf_2025), ensuring reproducibility.
3.4 Evaluation Protocol
Evaluation metric.
Building on the setup in §3.2, we now describe how in-context performance is evaluated. For each transformation $f\in F$, we draw an $n$-shot demonstration set $E\subset S$ and a held-out query $x\in S\setminus E$. The model $M$ receives a few-shot prompt of input–output pairs followed by the query and produces a prediction $\hat{y}=M(E,x)$. A trial is marked correct if $\hat{y}=f(x)$, giving the single-trial indicator $\mathcal{A}(f,E,x,M)=\mathbbm{1}\big[\hat{y}=f(x)\big].$
Since exact enumeration of all the programs is intractable, we use a Monte Carlo estimate (a full formulation of what we approximate can be found in the appendix B). Let $E_{n}$ denote the space of all sets of input bitstrings with cardinality $n$: $E_{n}=\{E\subset S:\,|E|=n\}$. Sample $m$ i.i.d. context sets $E^{(t)}\sim\mathrm{Unif}(E_{n})$ for $t=1,\ldots,m$; for each $t$, sample $x^{(t)}\sim\mathrm{Unif}(S\setminus E^{(t)})$. Then our empirical estimate of accuracy, for a given transformation $f$ is $\hat{\mathcal{A}}_{f}(M,n)\;=\;\frac{1}{m}\sum_{t=1}^{m}\mathcal{A}\big(f,\;E^{(t)},\;x^{(t)},\;M\big)$ and the overall estimated accuracy for model $M$ is:
$\hat{P}(M,n)=\frac{1}{|F|}\sum_{f\in F}\hat{\mathcal{A}}_{f}(M,n).$ (1)
Model suites and sampling configuration. We fix the number of Monte Carlo trials to $m=8$. The set of evaluated genomic models is $\mathcal{M}_{G}=\{\texttt{evo2\_1b\_base},\texttt{evo2\_7b},\texttt{evo2\_40b}\}$ and the set of linguistic models is $\mathcal{M}_{L}=\{\texttt{Qwen3-0.6B-Base},\texttt{Qwen3-1.7B-Base},\texttt{Qwen3-4B-Base},\texttt{Qwen3-8B-Base},\texttt{Qwen3-14B-Base}\}.$ We evaluate across shot counts $\mathcal{N}=\{1,2,4,8,16,32,64,128\}.$ For each model $M$ and shot count $n$, we compute:
$P_{L}=\{\hat{P}(M,n):M\in\mathcal{M}_{L},n\in\mathcal{N}\},\quad P_{G}=\{\hat{P}(M,n):M\in\mathcal{M}_{G},n\in\mathcal{N}\}.$
We estimate standard errors via a two-stage nonparametric cluster bootstrap, resampling functions (clusters) and within each selected function, resampling its evaluation samples with $5000$ replicates for each $(M,n)$.
Mode baseline. We define a mode baseline that always predicts the most frequent output observed in the context. For a given function $f$, context set $E\subset S$, and query input $x\in S\setminus E$, the mode prediction is $\hat{y}_{\text{mode}}(f,E,x)=\arg\max_{y\in S}\big|\{\,e\in E:f(e)=y\,\}\big|,$ where ties in the $\arg\max$ are broken randomly. This simply corresponds to guessing the most common output in the few-shot examples. The overall mode baseline with $n$ shots and across the set of all functions $F$ is:
$\hat{P}_{\text{mode}}(n)=\frac{1}{m|F|}\sum_{f\in F}\sum_{t=1}^{m}\mathbbm{1}\big[\hat{y}_{\text{mode}}(f,E^{(t)},x^{(t)})=f(x^{(t)})\big],$ (2)
where $E^{(t)}$ and $x^{(t)}$ are sampled identically to the model evaluation in Eq.1. This baseline corresponds to making an educated guess based only on the overall distribution of function outputs that simply learns the majority statistics of the prompt without attempting to infer the underlying transformation.
4 Empirical Results
This section reports empirical findings on few-shot bitstring generalization. §4.1 presents the main accuracy trends with respect to model size and shot count. §4.3 examines sensitivity to task complexity using a BitLoad measure, and §4.4 contrasts the models’ qualitative competencies across individual transformations.
4.1 Main Results
Figure 2: Few-shot performance of Qwen3 and Evo2 models. (a) Evo2 model performance with respect to log(shots). All models monotonically improve – the 7B and 40B have roughly equivalent performance, and the 1B trails behind them. (b) Qwen3 model performance with respect to log(shots). All models improve, but not always monotonically. Smaller models struggle in 4-16 shot range. (c) At comparable sizes, Evo2 outperforms Qwen3. (d) Averaged performance across both model families shows consistent improvement with respect to log(shots). All models exceed the mode baseline shown in gray color.
Across both model families, accuracy generally rises linearly with respect to $\log(\text{shots})$. A linear regression linking performance to $\log(\text{shots})$ yields highly significant slopes for all models (all $p\leq 10^{-3}$ via one-sided t-test on slope). As Fig. 2 shows, Evo2 models show cleanly monotonic gains, with a pronounced step from 1B to 7B, and near-indistinguishable curves for 7B and 40B; by 128 shots, both 7B and 40B surpass 40% accuracy (Fig. 2a). Qwen3 also trends upward overall but with non-monotonic patches—especially for smaller models in the 4–16 shot band—before resuming clear improvements from 32 to 128 shots; at 128 shots, the 14B model approaches 35% (Fig. 2b). A full table of all accuracies is in Appendix C.
Evo2 outperforms Qwen3 at comparable sizes. Evo2 1B beats Qwen3 1.7B and Evo2 7B beats Qwen3 8B at higher shot counts (Fig. 2c). Averaging within families reinforces the same picture: comparable performance in the 1–4 shot range, Evo2 pulling ahead around 8–16 shots, and the gap persisting at higher shot counts (Fig. 2d.)
The mode baseline’s performance doesn’t improve with more shots, and lags Qwen3 and Evo2 at high shot counts (Fig. 2d). This indicates that the in-context learning performed by Qwen3 and Evo2 is qualitatively different than simply sampling from the distribution of possible outputs in-context. Qwen3 and Evo2’s performance requires that they learn, in-context, how to condition on the input. Furthermore, for Qwen3, advantages over the mode baseline become consistently significant ($p<0.05$ via one-sided z-test on bootstrapped standard errors) at n=128n{=}128 for all sizes greater than 0.6B. For Evo2, statistically significant advantages emerge slightly earlier: for 1B at n=64n{=}64, the 7B at $n=32$, and the 40B at $n=16$. Regardless, all models surpass the naive baseline.
Figure 3: Performance at $n=128$ shots. All model accuracies increase monotonically with respect to parameter count.
4.2 Analysis of ICL Capability Across Model Families and Sizes
As shown in Fig. 3, both Qwen3 and Evo2 exhibit clear performance gains with increasing scale. Qwen3 displays a strong positive correlation between accuracy and parameter count ($p<0.05$ for all shot counts $\geq$16, one-sided $t$-test on slope), while Evo2 shows a distinct jump from 1B to 7B but little improvement beyond that. This suggests that in-context program induction becomes more robust with scale, though saturation may occur once models reach sufficient capacity. A detailed statistical analysis is provided in Appendix D.
4.3 ICL Sensitivity to Task Complexity: BitLoad Analysis
To understand which transformations Qwen3 and Evo2 infer most effectively, we analyze their performance across varying task complexities. We focus on the largest models in each family (Qwen3-14B and Evo2-40B) under the ($n=128$) shot regime, providing both models ample opportunity to display their ICL abilities.
Figure 4: Accuracy vs. BitLoad averaged across all tasks (BitLoad; Eq.3). Qwen declines sharply with increasing BitLoad, while Evo degrades more gradually, indicating greater robustness. Details in §4.3.
Defining BitLoad. We introduce BitLoad, a measure of a function’s intrinsic complexity. Informally, BitLoad quantifies how many input bits influence the output. Formally, it is defined as:
$\text{BitLoad}(f)=\sum_{i=1}^{k}\mathbbm{1}\Big[\exists x,j:f_{j}(x)\neq f_{j}(x^{\oplus i})\Big],$ (3)
where $x^{\oplus i}$ denotes $x$ with bit $i$ flipped, and $f_{j}(x)$ returns the $j$-th output bit. Intuitively, it counts the number of bit positions whose perturbation changes the output. So, the larger the BitLoad of a function, the harder it is since it requires the model to attend to more bits.
Fig. 4 shows the mean accuracy of Qwen vs Evo with respect to the BitLoad of all of our tasks. A full table of the BitLoad of every tested task is attached in the appendix A.3. We see that both models achieve near-perfect accuracy on constant (0 BitLoad) tasks and remain similar at BitLoad 1. However, by BitLoad 2, Evo begins to outperform Qwen, and beyond that point Qwen’s accuracy drops sharply – falling below 20% by BitLoad 4 – while Evo’s performance declines more gradually, remaining above 40% before converging near 20% at BitLoad 8. This asymmetry suggests that Evo maintains partial generalization as dependency depth increases (BitLoad value), whereas Qwen’s ICL collapses more abruptly.
While BitLoad strongly correlates with task accuracy, Fig. 4 also shows notable deviations within the error bars. Thus, while task complexity is a strong predictor of ICL accuracy, other factors (e.g., transformation depth, pretraining exposure) likely play meaningful roles, which motivate our qualitative study in §4.4.
4.4 Qualitative Analysis of Functional and Behavioral Differences
To complement the quantitative BitLoad analysis (§4.3), we conduct a per-task comparison to identify qualitative differences between Qwen3 and Evo2. Specifically, we analyze which tasks each model performs best on and how their inductive profiles diverge. We focus on the $n=128$ shot regime for the largest models – Qwen3-14B and Evo2-40B – where both have maximal opportunity to exhibit ICL. We rank all tasks by model accuracy and examine the top ten for each model. Tasks that appear in one model’s top ten but not the other are considered “exclusive.”
Exclusive competencies. Qwen’s exclusive tasks involve right-shift operations: "spread_last_bit" $\to$ "shift_right_zero" and "edge_mask" $\to$ "shift_right_zero". For instance, “shift_right_zero” pads the bitstring on the left with a zero and truncates the last bit (e.g., $01010000\to 00101000$). Qwen achieves 100% accuracy on "spread_last_bit" $\to$ "shift_right_zero", whereas Evo2 achieves only 50%. A similar but smaller gap appears for "edge_mask" $\to$ "shift_right_zero" (87.5% vs. 62.5%).
In contrast, Evo2’s exclusive strengths involve multi-bit transformations. A clear example is "flip_bits" $\to$ "right_half", which applies bitwise NOT followed by masking the first half of the input (e.g., $01011100\to 00000011$). Evo2 achieves 87.5% accuracy, while Qwen only 25%. This task has a BitLoad of four, consistent with Evo’s superior performance on medium-complexity (2–4 bit) transformations observed in Fig. 4.
Shared strengths. Despite these differences, 7 of the top 10 tasks are shared between Qwen and Evo, yielding an intersection-over-union of 0.54. Both models excel at simple transformations such as constant-outputs or single-bit dependencies, e.g., "spread_last_bit" which copies the final bit to all positions.
Differential skill profiles. To sharpen contrasts, we identify the ten tasks most favoring each model. Qwen’s advantages are concentrated in simple shifts and aggregation tasks. It outperforms Evo by 37.5% on the "minority" operation (output all 1s if zeros >> ones), and by a similar margin on two parity-based tasks requiring counting the number of 1s. These trends suggest that Qwen may be better at reasoning over global properties of bitstrings. Its superiority on simple shifts may also be explained by the extreme rarity of frame shift mutations in DNA due to how catastrophic they are – a single nucleotide offset can decimate an entire protein. This is empirically supported by the fact that single-nucleotide deletions/shifts increase perplexity far more than other common mutations when presented to Evo2 (evo2report).
Evo2, by contrast, dominates tasks requiring full-bitstring manipulation. It achieves 62.5% on bitwise NOT (vs. 0% for Qwen), 62.5% on identity (vs. 12.5%), and large margins on compositions such as "flip_bits" $\to$ "right_half" (87.5% vs. 25%) and "rotl1" $\to$ "flip_bits" (37.5% vs. 0%). This exposes perhaps the most important difference between Qwen and Evo’s ICL in this specific context: Evo can learn simple full-bitstring operations in-context, whereas Qwen cannot. Notably, Qwen3’s base models are trivially capable of learning the identity in a more familiar few-shot context – when examples are presented with arrows and newlines separating them, instead of our intentionally unfamiliar encoding. Thus these results should be taken as an existence proof of ICL in Evo2, not a definitive statement of Evo2 having more ICL ability than Qwen. We leave a comparison of these models across broader tasks to future work.
5 Discussion
What are the implications of our findings on prior efforts to explain the emergence of ICL? First, let us organize the existing frameworks for pinpointing the conditions under which ICL emerges:
  1. ($E_{1}$)
    ICL’s emergence is due to data distributional properties: The distributional properties of data, such as “parallel structures” in human language pretraining data (chen2024parallel), its compositional structure (hahn2023theory), “burstiness” (chan2022data) and other such properties (wibisono2024context; reddy2023mechanistic) may be of importance (and perhaps necessary) for the emergence of ICL.
  2. ($E_{2}$)
    ICL’s emergence is due to a compression mechanism: The large-scale compression mechanism during massive pretraining might drive ICL (elmoznino2024complexity; elmoznino2024context; hahn2023theory).
  3. ($E_{3}$)
    ICL’s emergence may require specific architectural properties: While Transformers might be better suited for ICL than LSTMs (xie2021explanation), evidence is mixed (lee2023exploring), and non-Transformer models have also demonstrated ICL capabilities (grazzi2024mamba; park2024can).
Our findings refine existing hypotheses about ICL’s origins. The emergence of ICL in genomic models challenges accounts that rely solely on language-specific distributional structures ($E_{1}$). The presence of ICL across both genomic and linguistic models supports the compression-based explanation ($E_{2}$), suggesting that large-scale sequence compression and its induced inductive biases drive ICL across modalities. With respect to architecture ($E_{3}$), Evo2 – an autoregressive hybrid combining convolutional and attention layers rather than a pure Transformer – exhibits similar scaling behavior, indicating that ICL does not depend on the pure Transformer form. Instead, architecture provides an expressive substrate that enables pattern induction once exposed to sufficiently large and structured data. Overall, these results position ICL as a modality-agnostic outcome of large-scale next-token prediction, rather than a phenomenon tied to linguistic statistics or a specific architecture.
What are the implications of our findings on the frameworks to explain how ICL operates? We next consider the major perspectives that seek to explain how ICL operates. One view holds that ICL functions as a mix of task learning and task retrieval, with demonstrations serving either to recall pretrained capabilities or to enable learning on the fly (pan2023incontext; lin2024dual; wang2024investigating; fang2025iclciphersquantifyinglearning). Our symbolic reasoning tasks, instantiated in both linguistic and genomic domains, provide direct evidence for this task learning mode, aligning with this hypothesis and prior work (pan2023incontext; fang2025iclciphersquantifyinglearning). Because these tasks do not depend on pretrained semantic priors, they do not invoke task retrieval, offering limited insight into the Bayesian view that interprets ICL as implicit inference over latent concepts (xie2021explanation; panwar2023context; wang2024large; jiang2024context). Meanwhile, our results remain agnostic toward the optimization-based hypothesis, which posits that ICL implements an implicit gradient-descent-like process (akyurek2022learning; ahn2023transformers; mahankali2023one; li2023closeness), as well as the induction-based account, which attributes ICL to specialized “circuits” for performing inductive generalization (elhage2021mathematical; olsson2022context; wang2023label; bansal2023rethinking; ren2024identifying). Together, our results most strongly support the presence of genuine task learning within ICL.
Could Evo2’s tokenizer grant it a unique advantage on these tasks? Evo2’s absolute advantage over Qwen3 may be explained by the fact that Evo2’s entire pretraining distribution consists of long-context reasoning over a set of only a few tokens, whereas this kind of in-context operation is far more alien to Qwen3. Further analysis would be needed to verify if this is what grants Evo2 its absolute advantage. Ideally, a model family trained on a third modality with a vocabulary between a few nucleotides and hundreds of thousands of words could be employed.
[…]
Beyond the H1-H2 dichotomy. While we presented H1 and H2 (in §1) as contrasting hypotheses, they are not necessarily mutually exclusive. It is entirely plausible that both hold simultaneously – that certain distributional properties inherent in natural data, shared across human language and other natural domains such as genomics, contribute to the emergence of ICL. In this view, linguistic compositionality may represent one instance of a broader statistical substrate that fosters ICL. We framed the two hypotheses as a dichotomy primarily to highlight the contrast between language-specific and modality-general explanations, rather than to suggest that only one can be true.
6 Conclusion
We introduce a suite of bitstring reasoning tasks that can be encoded in both natural language and genomic sequences, showing that genomic models – like their linguistic counterparts – exhibit clear in-context learning. Across all Evo2 model sizes, we observe robust log-linear gains in accuracy with increasing demonstrations, paralleling the scaling trends of Qwen3 language models. These findings challenge the notion that ICL is unique to human language, suggesting it emerges whenever an expressive model is trained autoregressively on structured, pattern-rich data.
Potential future work: This proof-of-existence for ICL lays the groundwork for many future research directions. One could broaden the test suite to incorporate tasks beyond bitstrings – up to four symbols are natively available in nucleotides, and far more can be used if one encodes at the codon level.
Another promising direction is mechanistic interpretability: analyzing Evo2’s internal activations to identify which circuits enable ICL and how the model aggregates context to predict the next nucleotide. Comparing these mechanisms to known induction circuits (elhage2021mathematical) in LLMs could reveal whether analogous structures exist across architectures – a question that connects biological and linguistic ICL dynamics.
Finally, this work motivates searching for ICL in other non-linguistic modalities – time series (das2024decoderonlyfoundationmodeltimeseries), system logs (akhauri2025performancepredictionlargesystems), physics simulations (pmlr-v267-holzschuh25a), chess games (ruoss2024amortizedplanninglargescaletransformers), and climate projections (duncan2025samudracefastaccuratecoupled). Each offers a structured, patterned substrate that could support its own form of contextual reasoning. These diverse modalities, each with their unique structure and constraints, suggest a rich world of non-linguistic ICL capability waiting to be explored, and this work represents a maiden voyage into these extremely interesting waters.
Broader Impact Statement
Understanding ICL is pivotal for both the scientific study and practical control of LLMs. ICL ties to major efforts to interpret model behavior, characterize how reasoning and abstraction evolve with scale, and design mechanisms for controllability and steering. It also powers pragmatic applications such as synthetic data generation. By demonstrating that ICL arises even in non-linguistic settings, this work broadens the empirical foundation for studying ICL as a general computational phenomenon rather than a quirk of human-language training. A clearer mechanistic understanding of how ICL emerges and operates across modalities can inform how we build, guide, and evaluate large models – improving their reliability, controllability, and usefulness while reducing risks from misalignment or overgeneralization.
Appendix A Further Details on Synthetic Task Definition
A.1 Table of Primitives
The following table describes the thirty primitives used to construct the task space via composition – their use is described in §3.2.
Primitive Description
alternating_start_one Produce a mask marking positions that differ from the alternating 1010… pattern starting with 1 (1 = mismatch, 0 = match).
alternating_start_zero Produce a mask marking positions that differ from the alternating 0101… pattern starting with 0 (1 = mismatch, 0 = match).
center_mask Zero out the first and last bits while leaving the interior bits unchanged; strings of length $\leq 2$ become all zeros.
double_rotl Circularly rotate the bitstring two positions to the left.
double_rotr Circularly rotate the bitstring two positions to the right.
edge_mask Preserve the first and last bits and zero out every interior bit (length 0/1 strings pass through).
flip_bits Invert every bit, swapping 0s for 1s and vice versa.
identity Return the bitstring unchanged.
invert_prefix Flip the bits in the left half of the string, keep the right half as is.
invert_suffix Keep the left half as is and flip every bit in the right half of the string.
keep_even_positions Keep bits at even indices (0-based) and zero out bits at odd indices.
keep_odd_positions Keep bits at odd indices (0-based) and zero out bits at even indices.
left_half Preserve the left half of the string and replace the right half with zeros.
majority Fill the string with the majority bit from the input; ties resolve to all 1s.
meta_constant Returns a random, pre-set constant.
minority Fill the string with the minority bit from the input; ties resolve to all 0s.
mirror_half Copy the left half of the string onto the right half in reverse order, keeping the center bit unchanged for odd lengths.
ones_if_palindrome Output all 1s if the input is a palindrome; otherwise output all 0s.
parity_fill Output all 1s when the input contains an odd number of 1s; otherwise output all 0s.
reverse_bits Reverse the order of the bits in the string.
right_half Zero out the left half and keep the right half unchanged.
rotl1 Circularly rotate the bitstring one position to the left.
rotr1 Circularly rotate the bitstring one position to the right.
shift_left_zero Shift the string left by one, dropping the first bit and appending a 0 on the right.
shift_right_zero Shift the string right by one, inserting a 0 on the left and dropping the last bit.
spread_first_bit Replace every position with the first bit of the input.
spread_last_bit Replace every position with the last bit of the input.
swap_halves Swap the left and right halves of the string.
swap_pairs Swap each adjacent pair of bits (positions 0/1, 2/3, …).
xor_with_s0 Can only be applied after another primitive. Computes the logical XOR between the original input $s_{0}$ and the output of the first primitive.
Table 1: The 30 unary primitives used to construct functions in $F$.
A.2 Table of Functions
The following table describes the one hundred specific compositions of primitives used to construct the evaluation suite in §3.2.
Function Function Function
identity spread_last_bit swap_halves $\to$ shift_left_zero
rotl1 invert_prefix swap_halves $\to$ shift_right_zero
reverse_bits invert_suffix shift_left_zero $\to$ swap_halves
flip_bits meta_constant shift_right_zero $\to$ swap_halves
swap_halves flip_bits $\to$ reverse_bits keep_even_positions $\to$ flip_bits
majority rotl1 $\to$ reverse_bits keep_odd_positions $\to$ flip_bits
minority reverse_bits $\to$ rotl1 flip_bits $\to$ keep_even_positions
parity_fill rotl1 $\to$ flip_bits flip_bits $\to$ keep_odd_positions
alternating_start_one swap_halves $\to$ reverse_bits edge_mask $\to$ flip_bits
alternating_start_zero swap_halves $\to$ flip_bits center_mask $\to$ flip_bits
left_half double_rotl $\to$ flip_bits shift_left_zero $\to$ keep_even_positions
right_half rotr1 $\to$ flip_bits shift_left_zero $\to$ keep_odd_positions
double_rotl spread_first_bit $\to$ flip_bits shift_right_zero $\to$ keep_even_positions
rotr1 spread_last_bit $\to$ flip_bits shift_right_zero $\to$ keep_odd_positions
double_rotr left_half $\to$ flip_bits keep_even_positions $\to$ reverse_bits
ones_if_palindrome right_half $\to$ flip_bits keep_odd_positions $\to$ reverse_bits
mirror_half flip_bits $\to$ left_half shift_left_zero $\to$ parity_fill
spread_first_bit flip_bits $\to$ right_half shift_right_zero $\to$ parity_fill
spread_last_bit double_rotl $\to$ reverse_bits parity_fill $\to$ shift_left_zero
invert_prefix rotl1 $\to$ swap_halves parity_fill $\to$ shift_right_zero
invert_suffix xor_with_s0 spread_first_bit $\to$ shift_left_zero
meta_constant flip_bits $\to$ xor_with_s0 spread_last_bit $\to$ shift_right_zero
shift_left_zero ones_if_palindrome $\to$ flip_bits spread_first_bit $\to$ keep_even_positions
shift_right_zero flip_bits $\to$ mirror_half spread_last_bit $\to$ keep_odd_positions
swap_pairs invert_prefix $\to$ reverse_bits spread_first_bit $\to$ edge_mask
keep_even_positions left_half $\to$ reverse_bits spread_last_bit $\to$ edge_mask
keep_odd_positions right_half $\to$ reverse_bits spread_first_bit $\to$ center_mask
edge_mask parity_fill $\to$ flip_bits spread_last_bit $\to$ center_mask
center_mask rotl1 $\to$ spread_first_bit rotl1 $\to$ shift_left_zero
xor_with_s0 shift_left_zero $\to$ flip_bits rotl1 $\to$ shift_right_zero
flip_bits $\to$ reverse_bits shift_right_zero $\to$ flip_bits shift_left_zero $\to$ rotl1
rotl1 $\to$ reverse_bits flip_bits $\to$ shift_left_zero shift_right_zero $\to$ rotl1
reverse_bits $\to$ rotl1 flip_bits $\to$ shift_right_zero reverse_bits $\to$ edge_mask
rotl1 $\to$ flip_bits swap_pairs $\to$ flip_bits reverse_bits $\to$ center_mask
swap_halves $\to$ reverse_bits shift_left_zero $\to$ reverse_bits edge_mask $\to$ shift_left_zero
swap_halves $\to$ flip_bits shift_right_zero $\to$ reverse_bits edge_mask $\to$ shift_right_zero
double_rotl $\to$ flip_bits spread_first_bit $\to$ flip_bits shift_left_zero $\to$ shift_left_zero
rotr1 $\to$ flip_bits shift_left_zero $\to$ edge_mask shift_left_zero $\to$ swap_pairs
Table 2: The complete set of 100 functions in $F$, consisting of 30 single primitives and 70 composed functions ($(f\to g)(x)=g(f(x))$).
A.3 BitLoad of Functions
Function / Composition BitLoad Function / Composition BitLoad
identity 8 rotl1 8
reverse_bits 8 flip_bits 8
swap_halves 8 majority 8
minority 8 parity_fill 8
alternating_start_one 8 alternating_start_zero 8
left_half 4 right_half 4
double_rotl 8 rotr1 8
double_rotr 8 ones_if_palindrome 8
mirror_half 4 spread_first_bit 1
spread_last_bit 1 invert_prefix 8
invert_suffix 8 meta_constant 0
flip_bits $\to$ reverse_bits 8 rotl1 $\to$ reverse_bits 8
reverse_bits $\to$ rotl1 8 rotl1 $\to$ flip_bits 8
swap_halves $\to$ reverse_bits 8 swap_halves $\to$ flip_bits 8
double_rotl $\to$ flip_bits 8 rotr1 $\to$ flip_bits 8
spread_first_bit $\to$ flip_bits 1 spread_last_bit $\to$ flip_bits 1
left_half $\to$ flip_bits 4 right_half $\to$ flip_bits 4
flip_bits $\to$ left_half 4 flip_bits $\to$ right_half 4
double_rotl $\to$ reverse_bits 8 rotl1 $\to$ swap_halves 8
xor_with_s0 0 flip_bits $\to$ xor_with_s0 0
ones_if_palindrome $\to$ flip_bits 8 flip_bits $\to$ mirror_half 4
invert_prefix $\to$ reverse_bits 8 left_half $\to$ reverse_bits 4
right_half $\to$ reverse_bits 4 parity_fill $\to$ flip_bits 8
rotl1 $\to$ spread_first_bit 1 shift_left_zero 7
shift_right_zero 7 swap_pairs 8
keep_even_positions 4 keep_odd_positions 4
edge_mask 2 center_mask 6
shift_left_zero $\to$ flip_bits 7 shift_right_zero $\to$ flip_bits 7
flip_bits $\to$ shift_left_zero 7 flip_bits $\to$ shift_right_zero 7
swap_pairs $\to$ flip_bits 8 shift_left_zero $\to$ reverse_bits 7
shift_right_zero $\to$ reverse_bits 7 swap_halves $\to$ shift_left_zero 7
swap_halves $\to$ shift_right_zero 7 shift_left_zero $\to$ swap_halves 7
shift_right_zero $\to$ swap_halves 7 keep_even_positions $\to$ flip_bits 4
keep_odd_positions $\to$ flip_bits 4 flip_bits $\to$ keep_even_positions 4
flip_bits $\to$ keep_odd_positions 4 edge_mask $\to$ flip_bits 2
center_mask $\to$ flip_bits 6 shift_left_zero $\to$ keep_even_positions 4
shift_left_zero $\to$ keep_odd_positions 3 shift_right_zero $\to$ keep_even_positions 3
shift_right_zero $\to$ keep_odd_positions 4 keep_even_positions $\to$ reverse_bits 4
keep_odd_positions $\to$ reverse_bits 4 shift_left_zero $\to$ parity_fill 7
shift_right_zero $\to$ parity_fill 7 parity_fill $\to$ shift_left_zero 8
parity_fill $\to$ shift_right_zero 8 spread_first_bit $\to$ shift_left_zero 1
spread_last_bit $\to$ shift_right_zero 1 spread_first_bit $\to$ keep_even_positions 1
spread_last_bit $\to$ keep_odd_positions 1 spread_first_bit $\to$ edge_mask 1
spread_last_bit $\to$ edge_mask 1 spread_first_bit $\to$ center_mask 1
spread_last_bit $\to$ center_mask 1 rotl1 $\to$ shift_left_zero 7
rotl1 $\to$ shift_right_zero 7 shift_left_zero $\to$ rotl1 7
shift_right_zero $\to$ rotl1 7 reverse_bits $\to$ edge_mask 2
reverse_bits $\to$ center_mask 6 edge_mask $\to$ shift_left_zero 1
edge_mask $\to$ shift_right_zero 1 shift_left_zero $\to$ shift_left_zero 6
shift_left_zero $\to$ swap_pairs 7 shift_left_zero $\to$ edge_mask 1
Table 3: BitLoad for every primitive and composed function in $F$. Used in §4.3.
Appendix B A Global Metric Over All Programs
For a given function $f$, model $M$, and number of in-context examples $n$, we define the average accuracy over all context sets $E_{N}=\{E\subset S:|E|=n\}$:
$A_{f}(M,n)=\frac{1}{|E_{N}|(|S|-n)}\sum_{E\in E_{N}}\sum_{x\in S\setminus E}A(f,E,x,M).$
The overall benchmark score across all functions $F$ is then
$P(M,n)=\frac{1}{|F|}\sum_{f\in F}A_{f}(M,n).$
Because exact evaluation is intractable (for $n=8$, $|E_{N}|(|S|-N)\;=\;\binom{|S|}{n}(|S|-n)\;=\;\binom{256}{8}\cdot 248\;\approx\;1.016\times 10^{17}$) we estimate $A_{f}(M,n)$ via Monte Carlo, as discussed in §3.4.
Appendix C Full Accuracy Results
Here we show the full table of accuracies used in §4.1 to analyze the ICL capabilities of Evo2 and Qwen3 and perform the necessary statistical tests.
Model 1 Shot 2 Shots 4 Shots 8 Shots 16 Shots 32 Shots 64 Shots 128 Shots
Qwen3 0.6B 13.4±2.3 14.1±2.3 16.0±2.6 19.5±2.7 16.1±2.4 16.8±2.5 19.1±2.7 23.6±2.9
Qwen3 1.7B 13.1±2.2 15.0±2.7 17.2±2.8 16.4±2.6 16.6±2.7 20.4±2.9 24.5±3.4 28.5±3.5
Qwen3 4B 13.5±2.4 15.2±2.4 17.0±2.6 15.4±2.5 18.4±2.7 24.8±3.2 24.5±3.4 29.4±3.4
Qwen3 8B 13.5±2.3 15.4±2.5 16.9±2.7 19.6±2.7 20.6±2.8 27.9±3.3 30.9±3.5 32.2±3.5
Qwen3 14B 14.0±2.4 14.1±2.4 16.8±2.7 20.6±2.9 22.9±3.1 25.5±3.2 32.6±3.7 33.8±3.5
Evo2 1B 13.1±2.2 15.0±2.4 17.1±2.9 18.2±2.7 22.0±2.9 23.8±2.8 26.5±3.0 30.8±3.1
Evo2 7B 14.5±2.4 15.8±2.7 17.6±2.9 22.6±2.9 25.2±3.0 29.8±3.1 32.2±3.1 41.0±3.4
Evo2 40B 15.0±2.6 15.5±2.5 17.6±2.8 21.6±3.1 26.4±3.1 28.9±3.1 34.5±3.2 41.1±3.3
Table 4: In-context learning performance across model families and shot counts. Values show accuracy ± standard error. All numbers are percentages. Bold numbers show the best performance within a model family. Models are ordered by parameter count within each family.
Appendix D Meta-Regression for ICL Efficacy with Number of Demonstrations
(a) Few-shot performance of Qwen3 and Evo2 models. All models show consistent linear improvement with respect to $\log(\text{shots})$. In contrast, no such improvement occurs for the naive baseline.
(b) Baseline accuracy decreases slightly with scale as the model gains more parameters for both Evo2 and Qwen3.
(c) ICL rate vs. model size: sharp gains up to  4B for Qwen3; mild boost from Evo2 1B to 7B; both plateau after 4–7B.
Figure 5: Few-shot behavior and scaling trends across Qwen3 and Evo2.
We perform linear regressions to predict accuracy from shot count with each model. For each model $M$, fit the linear regression: $\hat{P}(M,n)=\alpha_{0}(M)+\alpha_{1}(M)\log(n)+\varepsilon$. The raw regressions are shown in Fig. 5(a). Predictably, all $\alpha_{1}$ are positive as all models are capable of learning in-context.
We can interpret $\alpha_{0}$ as representing the model’s base accuracy at the task, what it would logically achieve with only one shot to identify the task. We can then interpret $\alpha_{1}$ as the model’s ICL efficacy: the speed at which it adapts to the task being presented and at which its accuracy improves. Analyzing how these values change across parameter values reveals insights into the ICL abilities of both the Qwen3 and Evo2 models.
First, we analyze how $\alpha_{0}$ changes in each model family – this analysis can be seen in Fig. 5(b). Both Evo2’s and Qwen3’s initial $\alpha_{0}$ remains essentially constant model-to-model, indicating that all models have similar levels of few-shot baseline performance. Notably, Evo2 and Qwen3 have essentially identical intercepts at around 0.12. This implies that despite drastically different training data, the overall amount of prior knowledge the models have coming into this task is roughly similar. This rules out Qwen simply having ‘less experience’ with this sort of task.
If one looks at $\alpha_{1}$ – ICL efficacy – in Fig. 5(c), a dramatically different picture is painted. Here, both Qwen3 and Evo2 follow similar patterns with a significant difference. Qwen3’s ICL efficacy increases monotonically with respect to parameters, more than doubling from the 0.6B to the 14B. Evo2 follows suit (albeit less dramatically), with ICL efficacy monotonically increasing from the 1B to the 40B.
In absolute terms, however, Evo2, when parameter-matched, adapts in-context faster than Qwen3 does. Evo2 40B outperforms Qwen3 14B significantly, and it takes until Qwen3 8B to exceed the ICL ability of Evo2 1B.
Taken together, this data suggests that Qwen3 and Evo2 have similar amounts of pretraining exposure to be able to solve these tasks, and that Evo2 simply has better overall ICL capability (in this regime) – even though Qwen’s ICL ability increases more rapidly with respect to parameters.
Appendix E Analysis of ICL with Program Synthesis Tasks: Bit-Diversity
BitLoad captures a measure of theoretical information bottleneck required to make predictions in program synthesis tasks. However, this does not always correlate with what models find easy to do. In Fig. 7 and Fig. 8, we plot the accuracy of Evo and Qwen models with respect to BitLoad. We notice that few high BitLoad tasks have high accuracy compared to all medium BitLoad tasks, illustrating that some high BitLoad tasks can be easy for these models to solve, probably due to the nature of patterns found during pre-training.
To further analyze the nature of ICL exhibited through these tasks, we define BitDiversity as the number of minority bits in the output string. In Fig. 9 and Fig. 10, we plot the accuracy with respect to BitDiversity. These plots try to estimate the effect of entropy in the output on model performance, and we see a more expected trend: models tend to perform better on low-entropy outputs, regardless of BitLoad. However, bin-wise performance trends are always increasing with shots, supporting the central hypothesis that ICL increases with increasing number of demos in these models.
Prevalence of 0-BitDiversity outputs.
Looking at the expected and predicted outputs of trials from our tasks (Fig. 11), we made a few interesting observations about 0-BitDiversity (BD) outputs.
  • Around 25% of true targets are 0-BD. This is a significantly high number as we only have 2 0-BD bit strings in all possible bit strings of any length $K$. It implies that many of our tasks create low BitDiversity outputs on random inputs, i.e., 0-BD outputs.
  • Models tend to produce a lot of 0-BD outputs, much higher than the number of 0-BD true targets in the low-shot regime. But this number quickly drops to the expected number with higher shots.
  • A large portion of the baseline (1-shot) performance can be explained by this prevalence of 0-BD outputs, but with more shots we get stronger evidence of ICL with increasing correctly predicted non 0-BD cases.
Understandable Mistakes. A potential confound is what we will call "understandable mistakes". These mistakes occur when the model outputs an incorrect answer that would be correct for some other programs given the few-shot context. Formally, a model’s output $y$ (see 3.4 for notation):
$y=M\Big(e_{1}\to f(e_{1}),e_{2}\to f(e_{2}),\dots,e_{N}\to f(e_{N}),x\Big)$
is an understandable mistake if $y\neq f(x)$ but there exists $f_{2}\in F$ such that $\forall e_{i}\quad f(e_{i})=f_{2}(e_{i})$ and $y=f_{2}(x)$.
These understandable mistakes are an alarming confound at low shot counts, but their effect vanishes by $N=16$. They occur most in tasks with low BitDiversity, which can often be confused with each other. Figure 6 below shows how understandable mistakes in Qwen3-4B’s inferences decay exponentially as the number of shots increases.
Figure 6: Qwen3-4B’s rate of understandable mistakes with respect to the number of shots. Despite starting at 16% in the one-shot regime, they fall to less than 1% by 8 shots and vanish entirely at 32 shots. This underscores how understandable mistakes are only a confound at very low shot counts and can essentially be ignored past 4 shots.
Noise due to Monte Carlo Trials. We use $m=8$ for our per-function Monte Carlo trials when we compute accuracy. This has little aggregate impact when comparing model performance across the entire suite, but drastically reduces the significance of results when comparing models and trends at the task-level. We only have eight discrete bands of accuracy at which we can estimate a model’s per-task performance – which reduces the expressiveness of regression. Worse, this can lead to models getting a lucky prompt or two with a 0-BD output, which raises performance to 12.5% or 25% without the model truly understanding the task. Addressing these confounds would simply require increasing $m$ by an order of magnitude or so. Alternatively, tasks of interest could be identified and $m$ selectively increased for those tasks to enable more nuanced analysis.
Figure 7: BitLoad vs Accuracy for all Evo models at all ICL shot-numbers. The per-BitLoad distribution of model performance at each shot level shows an uneven affinity of models for solving tasks at different BitLoads. However, all trends support our overall hypotheses.
Figure 8: BitLoad vs Accuracy for all Qwen models at all ICL shot-numbers.
Figure 9: BitDiversity vs Accuracy for all Evo models at all ICL shot-numbers. Model performance follows a more natural pattern of increasing performance with decreasing output entropy. This highlights that it is difficult for models to decipher ICL patterns for high entropy outputs.
Figure 10: BitDiversity vs Accuracy for all Qwen models at all ICL shot-numbers.
Figure 11: Enumerating cases of 0 BitDiversity. For our defined tasks, around 25% of the true targets have 0 BitDiversity (BD). With low-shots, models tend to produce a large number of 0-BD predictions. But this number decreases significantly with increasing shots and tends to match the actual prior value. Similarly, a majority of the baseline (1-shot) performance of models can be explained through 0-BD outputs. With a higher number of shots, the model starts learning higher entropy patterns and presents stronger evidence of ICL.